# supReferences

### v4

@@ 3.1 Almost sure behaviour of particle filter estimates
Let $N\geq 1$ and consider Algorithm 1 where ($\big{(}\{U_{t,2}^{n}\}_{n=1}^{N}\big{)}_{t\geq 1}$ is a sequence of independent sets of independent $\mathcal{U}(0,1)$ random variables. Then,
$\mathbb{P}\big{(}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\geq\kappa,\,\,i.o.\big{)}=\mathbb{P}\big{(}\|\eta_{t}^{N}-\eta_{t}\|\geq\kappa,\,\,i.o.\big{)}=1,\quad\forall N\geq 1,\quad\forall\kappa\in(0,1).$
3.2 Discussion
As shown in the above calculations, the negative result given in (6) (and thus in Proposition 2) for PFs (i.e. for Algorithm 1 where the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables) arises for the following simple and intuitive reason. On the one hand, since for all $t\geq 1$ the uniform random variables $\{U_{t,2}^{n}\}_{n=1}^{N}$ used by a PF are independent, each mutation step of the algorithm moves all the particle outside any given interval $[a,b]\subset\mathbb{R}$ with a constant positive probability. On the other hand, because the different mutation steps of a PF rely on independent sets of uniform random variables this event will necessarily happen infinity often.
Interestingly, in the literature it has been realized that if, in Algorithm 1, the function $G$ is defined by $G(x)=\mathbbm{1}_{\mathbb{R}\setminus I}(x)$ for some bounded interval $I\subset\mathbb{R}$ then, for a PF, it may happen that at some time $t\geq 1$ all the particles $\{X_{t}^{n}\}_{n=1}^{N}$ fall into $I$, in which case the particle system is said to be extinct (Del Moral,, 2004, see, Section 7.2.2, page 219). Surprisingly, it has not been realized that this positive probability to have, at any time $t\geq 1$, all the particles located in a given interval implies that PFs are not suitable for estimating an infinite sequence of filtering distributions.
Above we establish (6) using only the properties of the mutation steps of Algorithm 1. This observation has two key implications. Firstly, considering more sophisticated resampling schemes than those allowed by Algorithm 1 will not prevent (6) to hold. Secondly, this latter result remains true if for all $t\geq 1$ we replace the function $G(x)=e^{-cx^{2}/2}$ by some arbitrary function $G_{t}:\mathbb{R}\rightarrow(0,\infty)$. Notably, for the SSM considered in this work (6) holds for any sequence of observations $(y_{t})_{t\geq 1}$. More generally, by looking at the calculations used to prove (6) it seems hard to think of a meaningful SSM for which (6) does not hold. In particular, we stress that (6) is not due to the fact that we consider an unbounded state-space. To illustrate this point let $\mathsf{X}=[-l,l]$ for some constant $l\in(0,\infty)$, and let $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ be as defined in (3)-(4) where, for all $x^{\prime}\in\mathbb{R}$, $M(x^{\prime},\mathrm{d}x)$ denote the $\mathcal{N}_{1}(\rho x^{\prime},\sigma^{2})$ distribution truncated on $\mathsf{X}$ and where $\eta_{1}$ is such that $\eta_{1}(\mathsf{X})=1$. Then, by repeating the above calculations, it is trivial to verify that if in Algorithm 1 the $U_{t,2}^{n}$’s are independent $\mathcal{U}(0,1)$ random variables then
$\mathbb{P}\Big{(}\hat{\eta}_{t}^{N}\big{(}[-l,l]\setminus[-a,a]\big{)}=1,\,\,i.o.\Big{)}=\mathbb{P}\Big{(}\eta_{t}^{N}\big{(}[-l,l]\setminus[-a,a]\big{)}=1,\,\,i.o.\Big{)}=1,\quad\forall N\geq 1,\quad\forall a\in(0,l).$
3.3 Finite time horizon and probabilistic behaviour of particle filter estimates
As proved above, as time progress a PF will inevitably reach a time instant $\tau$ at which it will fail at estimating well the filtering and/or the predictive distribution of the SSM of interest. From a practical point of view it is of key interest to study the probability that such an undesirable event happens in a finite number $T$ of time steps.
Interestingly, Lemma 1, obtained for deterministic versions of Algorithm 1, can be used to address this problem. This is the case because if, in the algorithm, the $U_{t}^{n}$’s are independent $\mathcal{U}(0,1)^{2}$ random variables then $\sup_{t\in\{1,\dots,T\}}D^{*}_{2}(\{U_{t}^{n}\}_{n=1}^{N})\leq\sqrt{2/N}$ with high probability, provided that $N$ is sufficiently large. More precisely, using Aistleitner and Hofer, (2014, Theorem 1) we have the following key result:
Lemma 3.
Let $(s,T)\in\mathbb{N}^{2}$, $(\delta,q)\in(0,1)^{2}$ and, for all $N\geq 1$, let $\{\{U_{t}^{n}\}_{n=1}^{N}\}_{t=1}^{T}$ be $T$ independent sets of $N$ independent $\mathcal{U}(0,1)^{s}$ random variables. Then,
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}D^{*}_{s}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta\Big{)}\geq 1-q,\quad\forall N\geq N_{s,T,q}:=\frac{160s+33\log(T)-33\log(q)}{\delta^{2}}.$
By combining Lemma 1 and Lemma 3 we easily obtain the following result for PF with multinomial resampling:
Theorem 1.
Let $N\geq 1$ and consider Algorithm 1 where ($\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}_{t\geq 1}$ is a sequence of independent sets of $N$ independent $\mathcal{U}(0,1)^{2}$ random variables. Then, there exists a constant $\bar{C}_{1}\in[1,\infty)$ (independent of $N$) such that, for all $q\in(0,1)$ and all $T\in\mathbb{N}$, we have
$\mathbb{P}\Bigg{(}\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\bar{C}_{1}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T,q}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\,\,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\bar{C}_{1}\inf_{\gamma\in(0,1)}\frac{\delta_{N,T,q}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}}\Bigg{)}\geq 1-q$
with $\epsilon_{\star}\in(0,1)$ as in Proposition 1 and with $\delta_{N,T,q}=N^{-1}\big{(}160+33\log(T)-33\log(q)\big{)}$.
@@ 3.3 Finite time horizon and probabilistic behaviour of particle filter estimates
A direct application of Lemma 1 and of Lemma 3 would lead the conclusion of Theorem 1 to hold with $\delta_{N,T,q}$ replaced by $\delta_{N,T,q}^{1/2}$. In Theorem 1 we exploit a particular property of independent uniform random variables to improve this result (see the proof of the theorem for more details).
Remark 3.
In Theorem 1 the condition $C_{1}\geq 1$ comes from Lemma 3, and implies that the probabilistic bounds given in the theorem are meaningful (i.e. smaller than one) only when $N>N_{1,T,q}$, with $N_{1,T,q}$ as defined in Lemma 3.
A direct implication of Theorem 1 is that if, for some $\kappa\in(0,1)$, we have
$N\geq\Big{(}160+33\log(T)-33\log(q)\Big{)}\inf_{\gamma\in(0,1)}\kappa^{-\frac{2}{\gamma}}\bigg{(}\frac{\bar{C}_{1}}{1-\epsilon_{\star}^{1-\gamma}}\bigg{)}^{\frac{2}{\gamma}}$ (7)
then with probability at least $1-q$ we have $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ and $\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa$. This constitutes a reassuring result concerning the time evolution of PF estimates for two reasons. Firstly, it shows that, as the time horizon $T$ that we consider increases, the number of particles $N$ should grow at most at speed $\log(T)$ to ensure that the PF algorithm estimates the whole sets of distributions $\{\hat{\eta}_{t}\}_{t=1}^{T}$ and $\{\eta_{t}\}_{t=1}^{T}$ with a constant (probabilistic) error. Secondly, Theorem 1 establishes that $N$ should grow at most at speed $\log(1/q)$ if we want to increase the confidence level $1-q$.
The following proposition shows that these results concerning the dependence in $T$ and in $q$ of the number of particles needed to have both $\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa$ and $\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa$ with probability at least $1-q$ are sharp:
Proposition 3.
@@
Consider the set-up of Theorem 1. Then, for all $\kappa\in(0,1)$ there exist constants $\varrho_{\kappa}\in(0,1)$ and $T_{\kappa}\in\mathbb{N}$ such that, for all $q\in(0,1)$ and all $T\geq T_{\kappa}$, we have
$\mathbb{P}\Big{(}\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|\leq\kappa\Big{)}<1-q,\quad\forall N<\frac{\log(T)-\log(q)}{\log(1/\varrho_{\kappa})}.$
By using Lemma 2 and Theorem 1 we can easily obtain probabilistic bounds for the joint estimation of all the filtering and predictive expectations $\{\hat{\eta}_{t}(f)\}_{t=1}^{T}$ and $\{\eta_{t}(f)\}_{t=1}^{T}$ for a large class of functions, as illustrated with the following proposition:
Proposition 4.
Consider the set-up of Theorem 1 and let $\gamma\in(0,1)$ and $\mathfrak{F}\subset\mathcal{C}^{1}(\mathbb{R})$. Then, there exists a constant $C_{\mathfrak{F}}\in(0,\infty)$ such that, for all $T\geq 2$, all $q\in(0,1)$ and all $N\geq 1$, we have
$\mathbb{P}\bigg{(}\sup_{t\in\{2,\dots,T\}}\sup_{f\in\mathfrak{F}}\Big{(}\big{|}\hat{\eta}_{t}^{N}\big{(}f\mathbbm{1}_{[-a_{N},a_{N},]}\big{)}-\hat{\eta}_{t}(f)\big{|}\vee\big{|}\eta_{t}^{N}\big{(}f\mathbbm{1}_{[-a_{N},a_{N}]}\big{)}-\eta_{t}(f)\big{|}\Big{)}\leq C_{\mathfrak{F}}\,\delta_{N,T,q}^{\gamma\beta/2}\bigg{)}\geq 1-q$
for
  • $\mathfrak{F}=\big{\{}f\in\mathcal{C}^{1}(\mathbb{R}):\,\int_{\mathbb{R}}|f^{\prime}(x)|\mathrm{d}x=1\big{\}}$, if $a_{N}=\infty$ and if $\beta=1$.
  • $\mathfrak{F}=\big{\{}f\in\mathcal{C}^{1}(\mathbb{R}):\,|f(x)|\leq 1+|x|\text{ and }|f^{\prime}(x)|\leq 1,\,\forall x\in\mathbb{R}\big{\}}$, if $a_{N}=(1\wedge\delta_{N,T,q})^{-\gamma\alpha/2}$ for some constant $\alpha\in(0,1)$ and if $\beta=1-\alpha$.
  • $\mathfrak{F}=\big{\{}f\in\mathcal{C}^{1}(\mathbb{R}):\,|f(x)|\leq 1+e^{\zeta|x|}\text{and }|f^{\prime}(x)|\leq e^{\zeta|x|},\,\,\forall x\in\mathbb{R}\big{\}}$ for some constant $\zeta\in(0,\infty)$, if $a_{N}=-\gamma\alpha/(2\zeta)\log(1\wedge\delta_{N,T,q})$ for some constant $\alpha\in(0,1)$ and if $\beta=1-\alpha$.
In the last two examples given in Proposition 4, estimating e.g. $\hat{\eta}_{t}(f)$ with the truncated estimator $\hat{\eta}_{t}^{N}(f\mathbbm{1}_{[-a_{N},a_{N}]})$ is unusual (the typical estimate being the empirical counterpart $\hat{\eta}_{t}^{N}(f)$) but has an intuitive explanation. In these two examples the function $f$ is unbounded and thus, informally speaking, even a small estimation error of $\hat{\eta}_{t}$ in a region $R\subset\mathbb{R}$ where $\inf_{x\in R}|f(x)|$ is large will have a significant impact on the overall estimation error $|\hat{\eta}_{t}^{N}(f)-\hat{\eta}_{t}(f)|$.
4 Almost sure behaviour of sequential quasi-Monte Carlo
4.1 Remainder on scrambled $(t,s)$-sequences
Let $s\geq 1$ and $b\geq 2$ be two integers,
be the set of all $b$-ary boxes and $t\in\mathbb{N}_{0}$.
Then, for an integer $m\geq t$ the point set $\{u^{n}\}_{n=1}^{b^{m}}$ is called a $(t,m,s)$-net in base $b$ if every $b$-ary box of volume $b^{t-m}$ contains exactly $b^{t}$ points, while a sequence $(u^{n})_{n\geq 1}$ of points in $[0,1)^{s}$ is called a $(t,s)$-sequence in base $b$ if, for any integers $a\geq 0$ and $m\geq t$, the point set $\{u^{n}\}_{n=ab^{m}+1}^{(a+1)b^{m}}$ is a $(t,m,s)$-net in base $b$. From these definitions it should be clear that, from a practical perspective, $(t,s)$-sequences in base $b$ with $t=0$ and $b=2$ are preferable but such sequences exist only for $s\in\{1,2\}$, respectively (Niederreiter,, 1992, Corollary 4.24, page 62). In the context of this work, the crucial property of a $(t,s)$-sequence $(u^{n})_{n\geq 1}$ in base $b$ is to be such that, for some constant $C_{s,t,b}\in(0,\infty)$ depending only on $s$,$t$ and $b$, we have (see Niederreiter,, 1992, Section 4.1)
Finally, a sequence $(U^{n})_{n\geq 1}$ of $(0,1)^{s}$-valued random variables is a scrambled $(t,s)$-sequence in base $b$ if, $\mathbb{P}$-a.s., $(U^{n})_{n\geq 1}$ is a $(t,s)$-sequence in base $b$ and if $U^{n}\sim\mathcal{U}(0,1)^{s}$ for all $n\geq 1$. A scrambled $(t,s)$-sequence is obtained by randomizing, or scrambling, a $(t,s)$-sequence, and various scrambling methods have been proposed in the literature. The first $N$ points of a scrambled sequence can be generated in $\mathcal{O}(N\log(N))$ operations but the memory requirements of the algorithm needed to perform this task depend on the scrambling procedure that is used (Hong and Hickernell,, 2003). We refer to Chapter 3 in Dick and Pillichshammer, (2010) and to Owen, (2023) for a discussion on scrambling algorithms.
4.2 Main result
@@
In the context of this work we only need point sets in $(0,1)^{2}$ and thus, following the discussion of the previous subsection, we limit our attention to scrambled $(0,2)$-sequences in base $b=2$. More precisely, in what follows we study Algorithm 1 in the case where $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$. In this situation we have, using Niederreiter, (1992, Theorem 4.7, page 54 and Theorem 4.14, page 59),
$\sup_{t\geq 1}D^{*}_{2}\big{(}\{U_{t}^{n}\}_{n=1}^{N}\big{)}\leq\delta_{N},\quad\forall N\geq 1\quad\mathbb{P}-a.s.$
@@
A direct application of Lemma 1 then gives the following almost sure and time uniform guarantee for the resulting algorithm:
@@
Let $N\geq 1$ and consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$. Then, there exists a constant $\bar{C}_{2}\in(0,\infty)$ (independent of $N$) such that, with $\delta_{N}$ as defined in (9) and with $\epsilon_{\star}\in(0,1)$ as in Proposition 1,
$\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\bar{C}_{2}\inf_{\gamma\in(0,1)}\frac{\delta_{N}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\quad\sup_{t\geq 1}\|\eta_{t}^{N}-\eta_{t}\|\leq\bar{C}_{2}\inf_{\gamma\in(0,1)}\frac{\delta_{N}^{\gamma/2}}{1-\epsilon_{\star}^{1-\gamma}},\quad\mathbb{P}-a.s.$
@@
By using Lemma 2 and Theorem 2 we can easily obtain almost sure time uniform bounds for the SQMC estimates of the filtering and predictive expectations $(\hat{\eta}_{t}(f))_{t\geq 1}$ and $(\eta_{t}(f))_{t\geq 1}$ for a large class of functions. For instance, for SQMC the conclusion of Proposition 4 holds with $T=\infty$, $q=1$ and with $\delta_{N,T,q}$ replaced by $\delta_{N}$.
4.3 Discussion

### v5

@@ 3.2 Finite time horizon and probabilistic behaviour of particle filters
Consider the set-up of Theorem 1 and let $\bar{q}\in(0,1/2)$. Then, for all $\kappa\in(0,1]$, there exist constants $C_{\kappa}\in(0,\infty)$ and $T_{\kappa}\in\mathbb{N}$ such that, for all $q\in(0,\bar{q}]$ and all $T\geq T_{\kappa}$, we have
$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\leq\kappa\Big)<1-q,\quad\forall N<C_{\kappa}\log(T/q).$
@@
Using (2), it follows that if we replace the discrepancy metric $\|\cdot\|_{\mathrm{D}}$ by the Kolmogorov metric $\|\cdot\|$ then the conclusion of Proposition 2 holds only for all $\kappa\in(0,1/2)$.
3.2.1 Proof of Proposition 3
Proof of Proposition 3.
@@ 3.2.1 Proof of Proposition 3
Let $\kappa\in(0,1)$ be arbitrary and remark that, by Proposition 1 and (2), there exist constants $a_{\kappa}\in(0,\infty)$ and $T_{\kappa}\in\mathbb{N}$ such that, for all $t\geq T_{\kappa}$, we have both $\eta_{t}([-a_{\kappa},a_{\kappa}])>\kappa$ and $\hat{\eta}_{t}([-a_{\kappa},a_{\kappa}])>\kappa$. In addition, let $\varrho_{\kappa}=\min\big\{1-\Phi(4a_{\kappa}/\sigma),1/2\big\}$ and, for all $N\geq 1$, let $(\mathcal{F}^{N}_{t})_{t\geq 0}$ be as defined in the proof of Proposition 2. Then, by using (8), we have
$\mathbb{P}\Big(\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}^{N}_{t-1}\Big)\geq\varrho^{N}_{\kappa},\quad\forall t\geq 1,\quad\forall N\geq 1.$ (9)
Therefore, noting that for all $t\geq T_{\kappa}$ and $N\geq 1$ we have
$1-\mathbb{P}\Big(\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\leq\kappa,$ $\,\,\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\leq\kappa|\mathcal{F}_{t-1}^{N}\Big)\geq\mathbb{P}\Big(\hat{\eta}_{t}^{N}([-a_{\kappa},a_{\kappa}])=0,\,\,\eta_{t}^{N}([-a_{\kappa},a_{\kappa}])=0|\mathcal{F}_{t-1}^{N}\Big)$
it follows from (9) that, for all $T\geq T_{\kappa}$ and $N\geq 1$,
$\mathbb{P}\Big(\sup_{t\in\{1,\dots,T\}}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|_{\mathrm{D}}\leq\kappa,\,\,\sup_{t\in\{1,\dots,T\}}\|\eta_{t}^{N}-\eta_{t}\|_{\mathrm{D}}\leq\kappa\Big)\leq(1-\varrho^{N}_{\kappa})^{T-T_{\kappa}+1}.$ (10)
To conclude the proof of the proposition we let $T\geq T_{\kappa}$, $\bar{q}\in(0,1/2)$ and $q\in(0,\bar{q}]$ be fixed, and to simplify the notation we let $v_{T}=T-T_{\kappa}+1$. Then, we first note that
$(1-\varrho^{N}_{\kappa})^{v_{T}}<1-q,\quad\forall N<N_{T,\kappa,q}:=\frac{-\log\big(1-(1-q)^{1/v_{T}}\big)}{\log(1/\varrho_{\kappa})}$ (11)
where, using Bernoulli’s inequality (see e.g. Li and Yeh,, 2013, Theorem 3.1),
$(1-q)^{1/v_{T}}=\frac{(1-q)^{1+1/v_{T}}}{1-q}\geq\frac{1-q-q/v_{T}}{1-q}.$
The latter inequality implies that $1-(1-q)^{1/v_{T}}\leq(q/v_{T})/(1-q)\leq(q/v_{T})/(1-\bar{q})$, and thus
$N_{T,\kappa,q}\geq\frac{\log(v_{T}/q)+\log(1-\bar{q})}{\log(1/\varrho_{\kappa})}\geq\frac{\log(v_{T}/q)}{\log(1/\varrho_{\kappa})}\delta_{\bar{q}},\quad\delta_{\bar{q}}:=1+\frac{\log(1-\bar{q})}{\log(1/\bar{q})}$ (12)
where $\delta_{\bar{q}}>0$ since $\bar{q}\in(0,1/2)$. The result of the proposition follows from (10)-(12), upon noting that if $T_{\kappa}>1$ then
$\frac{\log(v_{T}/q)}{\log(T/q)}\geq\frac{\log(v_{T})}{\log(T)}\geq\inf_{x\in[T_{\kappa},\infty)}\frac{\log(x-T_{\kappa}+1)}{\log(x)}>0.$
where the first inequality uses the fact that $T\geq v_{T}$.
3.2.2 Some remarks
Corollary 1 constitutes a reassuring result concerning the time evolution of PF estimates. Indeed, it notably shows that, as the time horizon $T$ that we consider increases, the number of particles $N$ only needs to grow at speed $\log(T)$ to ensure that the PF algorithm estimates the whole sets of distributions $\{\hat{\eta}_{t}\}_{t=1}^{T}$ and $\{\eta_{t}\}_{t=1}^{T}$ with a constant (probabilistic) error.
It is important to stress that the result of Proposition 3 is not specific to the considered filtering problem but will hold in general. Indeed, by inspecting the calculations of Section 3.2.1, it should be clear that the conclusion of Proposition 3 holds as soon as the result of Proposition 1 and (7) hold. The former result is standard while, as we argued in Section 3.1.2, the latter result is expected to hold for a large class of filtering problems.
3.3 Almost sure behaviour of sequential quasi-Monte Carlo
@@
We now turn to the study of SQMC. In the context of this work we only need point sets in $(0,1)^{2}$ and thus, following the discussion in Section 1.4, we limit our attention to scrambled $(0,2)$-sequences in base $b=2$. More precisely, in this subsection we study Algorithm 1 in the case where $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$, in which case we have, by Niederreiter, (1992, Theorem 4.7, page 54 and Theorem 4.14, page 59),
$\sup_{t\geq 1}D^{*}_{2}\big(\{U_{t}^{n}\}_{n=1}^{N}\big)\leq\delta_{N},\quad\forall N\geq 1\quad\mathbb{P}-a.s.$ (13)
@@
For the resulting filtering algorithm we have the following almost sure and time uniform guarantee:
@@
Let $N\geq 1$ and consider Algorithm 1 where $\{U_{t}^{n}\}_{n=1}^{N}$ is the first $N$ points of a scrambled $(0,2)$-sequence in base $b=2$ for all $t\geq 1$. Then, there exists a constant $\bar{C}_{2}\in(0,\infty)$ (independent of $N$) such that, with $\delta_{N}$ as defined in (14),
$\max\bigg\{\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|,\,\,\sup_{t\geq 1}\|\eta_{t}^{N}-\eta_{t}\|\bigg\}\leq\bar{C}_{2}\,\delta_{N}^{1/2}\log\big(1+\delta_{N}^{-1/2}\big),\quad\mathbb{P}-a.s.$
@@
Theorem 2 implies that the SQMC estimates $(\hat{\eta}_{t}^{N})_{t\geq 1}$ and $(\eta_{t}^{N})_{t\geq 1}$ of the two sequences $(\hat{\eta}_{t})_{t\geq 1}$ and $(\eta_{t})_{t\geq 1}$ are such that $\lim_{N\rightarrow\infty}\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|=\sup_{t\geq 1}\|\eta_{t}^{N}-\eta_{t}\|=0$ with $\mathbb{P}$-probability one. Moreover, it allows to readily obtain the following result:
Corollary 2.
Consider the set-up of Theorem 2. Then, there exists a constant $\bar{C}^{\prime}_{2}\in(0,\infty)$ such that, for any $(q,\kappa)\in(0,1)^{2}$ and $T\geq 1$, the following implication holds:
$\frac{N}{\log(1+N)^{2}}\geq\bar{C}_{2}^{\prime}\big\{1+$ $\log(1+\kappa^{-1})\big\}^{2}\kappa^{-2}\implies\mathbb{P}\Big(\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa,\,\,\sup_{t\geq 1}\|\hat{\eta}_{t}^{N}-\hat{\eta}_{t}\|\leq\kappa\Big)=1.$
3.3.1 Some remarks